Candice paid her library fine with 10 coins consisting of nickels and dimes. If the fine was 80 cents, then how many of each type of coin did she use?
Answer by richwmiller(17219)

n+d=10
5n+10d=80
d=6
n=4
60 +20=80
Reminder:
You are limited to one problem per submission.
No more than 4 submissions daily
No similar problems.